OpenStudy (mellisa):

Solve the equation. Note if the equation is an identity or if it has no solution. -3(-c-18)=3c-18 A.c=___ B. The equation is an identity C. the equation has no solution

OpenStudy (danjs):

did you try and solve that to get a value for c?

OpenStudy (mellisa):

is it -12?

OpenStudy (danjs):

no, you can't solve that for c. -3(-c-18)=3c-18 (-3)*-c - 18*(-3) = 3c - 18 3c + 54 = 3c - 18 54 = -18

OpenStudy (mellisa):

ok thanks, can u help with a few others

OpenStudy (danjs):

That is a contradiction, a false statement. The lines are parallel, and have no solution where they intersect

OpenStudy (danjs):

Sure

OpenStudy (mellisa):

Use a table to find the solution of the equation. If the solution lies between two consecutive​ integers, identify those integers. 7.2=3n-4 If the solution is an​ integer, type this integer.​ Otherwise, type the two consecutive integers between which the solution lies.

OpenStudy (danjs):

solving for n. 7.2 = 3n - 4 11.2 = 3n n = 11.2 / 3

OpenStudy (danjs):

That is not an integer answer. 11.2/3 is about 3.7333333 that lies between integers 3 and 4

OpenStudy (mellisa):

cool ok

OpenStudy (mellisa):

Tell whether the given ordered pair is a solution. y=5-3x. (-1,-8) Does the ordered pair -1,-8 satify the solution? A.yes Bno

OpenStudy (danjs):

(x , y) = ( -1 , -8) Just have to put those values, x=-1, y=-8, into the equation and see if it works... y = 5 - 3x -8 = 5 - 3*(-1)

OpenStudy (mellisa):

so it is A. yes

OpenStudy (danjs):

-8 = 5 - (-3) -8 = 5 + 3 -8 = 8 that is not true

OpenStudy (mellisa):

oh so B.no sorry im kinda bad with math, atleast this stuff

OpenStudy (danjs):

no problem, have to practice it

OpenStudy (mellisa):

ok thanks that was my last one, if i need u again ill tag u

OpenStudy (danjs):

welcome